Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used in production (not on the basis of raw materials purchased). Its predetermined overhead rate was based on a cost formula that estimated $124,600 of manufacturing overhead for an estimated allocation base of $89,000 direct material dollars to be used in production. The company has provided the following data for the just completed year:

Purchase of raw materials $ 131,000
Direct labor cost $ 85,000
Manufacturing overhead costs:
Indirect labor $ 109,400
Property taxes $ 8,000
Depreciation of equipment $ 18,000
Maintenance $ 15,000
Insurance $ 10,600
Rent, building $ 37,000
Beginning Ending
Raw Materials $ 25,000 $ 18,000
Work in Process $ 49,000 $ 36,000
Finished Goods $ 70,000 $ 56,000

Required:

1. Compute the predetermined overhead rate for the year.

2. Compute the amount of underapplied or overapplied overhead for the year.

3. Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials.

4. Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer.

5. Assume that the $36,000 ending balance in Work in Process includes $8,400 of direct materials. Given this assumption, supply the information missing below:
Answer #1

1) Predetermine overhead rate = 124600/89000 = 140% of direct material used

2) Applied overhead = (25000+131000-18000)*1.4 = 193200

Actual overhead = 198000

Under applied overhead = 198000-193200 = 4800

3) Schedule of cost of goods manufactured

Direct material
Beginning material 25000
Material purchase 131000
Material available for use 156000
Less: Ending material -18000
Direct material used 138000
Direct labor 85000
Manufacturing overhead 193200
Total manufacturing cost 416200
Beginning work in process 49000
Total cost of work in process 465200
Less; Ending work in process -36000
Cost of goods manufactured 429200

4) Unadjusted cost of goods sold = 70000+429200-56000 = 443200

5) Calculate following

Direct material 8400
Direct labor 15840
Manufacturing overhead (8400*1.4) 11760
Ending WIP 36000